Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

If a person purchases 15 of the 3,000 tickets sold in a...

Expert replies
by swerve » Wed Jan 31, 2018 5:33 am
If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that this person will not win?

A. 0
B. 1/200
C. 1/2
D. 199/200
E. 1

The OA is D.

Please, can any expert explain this PS question for me? I tried to solve it but I can't get the correct answer.

I know that the probability that this person will win is 15/3000, but how can I determine the probability that this person will not win?. I need your help. Thanks.
Join the discussion
Source: — Problem Solving |

by DrMaths » Wed Jan 31, 2018 7:50 am
If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that this person will not win?

A. 0
B. 1/200
C. 1/2
D. 199/200
E. 1

Two solutions, same answer:

(1) Pure method:

P = P(winning ) = 15/3000 = 5/1000 = 1/200
P' = (P of not winning) = 1 - P = 1 - 1/200 = 199/200

(2) Looking at the given answers:

A. 0 - It is not impossible to lose!
B. 1/200 - A disproportionately small value!
C. 1/2 - It's clearly not an even chance!
D. 199/200 - is a reasonable figure
E. 1 - It is not certain to lose!
Join the discussion

by EconomistGMATTutor » Wed Jan 31, 2018 11:10 am
If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that this person will not win?

A. 0
B. 1/200
C. 1/2
D. 199/200
E. 1

The OA is D.

Please, can any expert explain this PS question for me? I tried to solve it but I can't get the correct answer.

I know that the probability that this person will win is 15/3000, but how can I determine the probability that this person will not win?. I need your help. Thanks.
Hi swerve,
Let's take a look at your question.

We know that we can find the probability of an event by dividing the number of successful outcomes divided by the total number of outcomes.
Number of tickets purchased by a person = 15
Total number of tickets = 3000

We will first find the probability of winning.
$$P\left(\text{The person will win}\right)=\frac{15}{3000}=\frac{1}{200}$$

Now, we can find the probability that the person will not win just by subtracting the above probability from 1.
$$P\left(\text{The person will not Win}\right)=1-\frac{1}{200}=\frac{200-1}{200}=\frac{199}{200}$$

Therefore, Option D is correct.

Hope it helps.
I am available if you'd like any follow up.
GMAT Prep From The Economist
We offer 70+ point score improvement money back guarantee.
Our average student improves 98 points.

Image
Join the discussion

by Scott@TargetTestPrep » Fri Feb 02, 2018 11:13 am
swerve wrote:If a person purchases 15 of the 3,000 tickets sold in a raffle that awards one prize, what is the probability that this person will not win?

A. 0
B. 1/200
C. 1/2
D. 199/200
E. 1
The probability of not winning is (3,000 - 15)/3,000 = 2,985/3,000 = 597/600 = 199/200.

Alternate solution:

The probability of winning is 15/3,000 = 1/200; thus the probability of not winning is 1 - 1/200 = 199/200.

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion